OpenStudy (anonymous):

[5(x - 4) - 2] + [5(x - 1) + 5]

OpenStudy (anonymous):

(5x - 20 - 2) + (5x - 5 + 5) =(5x - 22) + (5x) =10x - 22

jimthompson5910 (jim_thompson5910):

[5(x - 4) - 2] + [5(x - 1) + 5] [5(x)+5(-4) - 2] + [5(x)+5(-1) + 5] [5x-20 - 2] + [5x-5 + 5] [5x-22] + [5x] 5x-22 + 5x (5x+5x)-22 10x-22 So [5(x - 4) - 2] + [5(x - 1) + 5] = 10x-22

OpenStudy (anonymous):

the answer Jim gave is correct.